# Shahriar Noroozizadeh

### v2

@@ 4.3 Where does the low-rank spectral bias come from?
Open Question 2: Characterizing and Controlling the Geometric Bias (i) To what precise extent does a deep sequence model memorize different parts of a graph dataset geometrically vs. associatively? (ii) What local or global notions of graph complexity—such as the graph spectrum, connectivity, node & vertex count and degree—and what aspects of optimization—such as the architecture, learning rate, training time —control this bias? (iii) What interventions in the data, architecture or optimization can be made to make memorization more geometric or more associative as desired in practice?
4.4 The importance of weight-tying
Our Node2Vec model shares weights between the two layers, reflecting the de facto practice of tying unembedding and embedding matrices in language models. This convention coincidentally plays a unique role in memory: without it, the geometries do not seem visible, at least on the face of it (see Fig.˜10 for Node2Vec and Fig.˜11 for Transformer, and the corresponding multi-hop cosine-similarities in Fig.˜33 and Fig.˜34).
We give an explanation for this. A weight-untied two-layer model must factorize the adjacency matrix not as a eigenvalue decomposition (EVD) but as a singular value decomposition (SVD)—this intuitively follows from analyses of weight-untied Word2Vec-style models [199, 83, 152]. The eigenvalues of the adjacency matrix range both positive and negative; the SVD “folds up” the negative eigendirections—all strong eigendirections, both the most positive and most negative, become top singular directions (see [197] for relevant facts on spectral graph theory). The most negative eigendirections however assign zig-zagging values to adjacent nodes (e.g., +1+1 and 1-1 as we walk along a path). This manifests as the unruly figures in Figs.˜10 and 11.
@@ 4.4 The importance of weight-tying
This raises various questions. First, although visually unruly, these zig-zagging eigendirections are highly-structured: they approximate a bipartite cut of the graph. Thus, they correspond to a form of geometric memory distinct from the “global” one, best termed as a bipartite geometric memory. Next, while the “global” geometry endows the model with navigational powers (as witnessed in §2), what benefits does the bipartite memory bring to a model’s reasoning abilities?
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 10: Embeddings of weight-untied Node2Vec do not show a clean geometry in the top directions. Corresponding multi-hop cosine similarities are in Fig.˜33.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 11: Embeddings of a weight-untied Transformer do not show a clean geometry in the top directions. Corresponding multi-hop cosine similarities are in Fig.˜34.
@@ 5.1 Analysis of Transformer memory
Associative memory. The concept of associative memory dates back to theories of how information is stored in the brain [101, 182]. These ideas have since been explicitly modeled through various architectures such as Hopfield networks [70, 142], energy-based models [89], and other modern Transformer-style inventions [91, 69]. Closest to us are works that analyze how architectures implicitly behave as associative memory storage, such as in autoencoders [139] or Transformers [20, 26, 124, 49, 154, 164]. Others have established connections between the attention mechanism and Hopfield networks [142, 161]. We emphasize that this view is sufficient to understand Transformer behavior on disjoint facts, evidenced by the rich empirical literature built on this view. The geometric view only seems necessary when the facts become interdependent.
@@
Geometric memory as node representations. Although, geometric memory has not been explicitly labeled as a distinct form of parametric memory, we find scattered observations of it in literature in weaker forms. [78, 85, 189] show that nodes in graph-like datasets are embedded in a geometric way that encodes distances. These settings involve some form of global supervision or latent high-level structure, while we find that such a geometry can arise even with atomic facts. Jiang et al. [78] further juxtapose these structured embeddings against associative memory. However, they suggest that the structure becomes relevant only in the underparameterized setting, and that an associative memory suffices as an abstraction in the overparameterized setting; we find this is not the case. Independently, on two-hop datasets, [189, 174] report how models first “memorize” the training data, before “generalizing” on their reasoning task; under the hood, this translates to associative and geometric memorization respectively.
@@
Geometric memory as factorized knowledge. Recall that the geometric memory can be viewed as storing a low-rank factorization of the adjacency matrix. The advantages of such factorizations is again scattered in literature. Saxe et al. [153, 152] prove that a two-layer model naturally learns an SVD factorization of input-output associations. This is a powerful result demonstrating the value of depth in learning hierarchical representations; but the global, multi-hop nature of these representations was not highlighted. Also, unlike our setting, theirs requires explicit pressures to demonstrate a low rank spectral bias (either through a bottleneck or early-stopping) for they study the squared-error loss. [^10] Under their squared error loss (applied without a softmax layer), a fully-trained 2-layered model would learn the 1-hop adjacency matrix perfectly, with no global, multi-hop information revealed in the logits. Although the associations are stored in a factorized form, the final hidden representations do not filter out the lower directions. More recently, Huang et al. [74] point out that two-hop reasoning emerges if and only if the key-query matrices in an attention layer are kept separate (as against being collapsed into a single matrix); we show that such favorable effects occur in more general architectures without attention layers. They also highlight the role of an implicit regularization effect from their reasoning supervision; we emphasize that this does not resolve our memorization puzzle—see Remark˜1.
@@
We also interpret these prior analyses [153, 74] as a study of how $2$-layered models memorize a special type of graphs: bipartite graphs (although they do not express it as such). [^11] Concretely, [153] study mappings between “entity” tokens and binary feature vectors; we translate this as a bipartite graph where “entity” nodes are connected to various “feature” nodes. [74] study a mapping between “subject” nodes and “answer” nodes. Importantly, as noted in our own analysis in §4, such 2-layered models do not capture the competition between associative and geometric memory. [^12] At least not in its full glory; in $2$-layered weight-tied models, there is only a contrived form of associative memory (Proposition 5); if weight-untied, a natural form can arise in a single step (Proposition 4). For that, a study of deeper models is required, where the dynamics along the various spectral components become entangled; to keep them untangled, unnatural assumptions about the initializations need to be made, as shown in [152]. Another nuance in these prior 2-layered analyses is the lack of weight-tying, which creates significant differences in what is learned, as we identify in §4.4.
@@
(Implicit regularization in matrix factorization is an insufficient explanation of geometric memory) Huang et al. [74] attribute the success of reasoning to implicit regularization effects that arise in matrix completion tasks; we argue that these effects do not explain why a geometry arises over associative storage in our experiments. The “matrix completion” view in [74] is that during training the model attempts to fit certain observed entries in a $2$-hop adjacency matrix ${\bf{A}}^{2}$; any test query in effect demands that the model “complete” a missing $2$-hop entry. While many minima exist in such underdetermined matrix completion tasks, it is known [62, 11] that gradient descent on a deep factorized model finds the least nuclear norm matrix consistent with the observed entries. However, our edge-memorization-only task is a well-determined task that requires storing the adjacency matrix $\bf{A}$} as is, without any missing entries to infer; thus no such implicit pressure must exist. Yet a geometry arises in this task as shown in Section˜3.1. While it is still possible that there is a certain nuclear norm minimization effect even in edge-memorization tasks—perhaps because the logits themselves are not well-determined—the connection is not as obvious as it may seem.
@@
Expressive capacity. Theoretical works have quantified bounds on the expressive capacity of models when it comes to memorizing sequences [106, 81, 105, 80, 87] as opposed to associations between pairs of bigrams. These do not comment on the learning dynamics. These works typically assume that the token embeddings are all well-separated from each other (an assumption that empirically breaks in our setting). In light of the geometric view, it is necessary to restate expressive capacity bounds in terms of geometric capacity of a network and the “geometric complexity” of the dataset.
@@
Empirical analyses. Other works [10, 115, 144, 128] have performed careful empirical analyses of scaling laws for memorization, and quantified memorization in terms of “bits per parameter count”, known as bit complexity [171], which is related to our notion of bit count in Section˜3.2. All these analyses are worth revisiting with the newfound lens of a geometric memory. Zucchet et al. [203] empirically analyze the dynamics behind how facts are memorized in a model. Others [99, 196] have proposed methodological improvements to acquiring knowledge in a Transformer; of relevance to us is the finding in Zhang et al. [196] that training a model simultaneously on both facts and question-answering is a better way to integrate knowledge into the parameters.
@@
Mechanistic interpretability. There have been mechanistic investigations into how Transformers perform fact recall [104, 50] and where facts are stored in a transformer [49] and how it can be edited [201, 111]. Similar attempts have been made in traditional classifier networks [15, 107, 163]. Directly related to us are the works of Khona et al. [85] and Yao et al. [188], Biran et al. [21] who perform a mechanistic interpretability analysis of how multi-hop recall works.
@@ 5.2 In-weights reasoning tasks
Synthetic graph tasks. Our work consolidates positive results of in-weights reasoning in literature, and presents a stronger instance of it, removing various confounders that make composition-learning easy. Khona et al. [85] report successful path-finding on $200$ nodes-large in-weights graphs, with varying path lengths and test-train overlap. Others [174, 43, 189, 74] report positive results on much shorter 2-hop tasks over a thousand entities or fewer. Geerts et al. [48] look at in-weights transitive inference, a special type of $\ell$-fold composition query where the model is trained on local comparisons and is queried on more distant comparisons. On settings with $7$ objects, Geerts et al. [48] find a clear difference between an in-context version of their task (where the model struggles) and an in-weights version (where the model succeeds). Our task of finding the first node is a much harder search task akin to finding the smallest node in an ordered relationship. Nagarajan et al. [119] discuss the limitations of next-token prediction, including on open-ended in-weights tasks, in lower data regimes. The fact that their next-token predictor achieves non-trivial performance on their in-weights task could be attributed to the effects of a geometric memory. Tangentially related is the positive finding in Yin and Wang [191] that the Transformer can compose in-weights knowledge given in-context demonstrations. It is worth noting that the (theoretical) arguments in both these works [119, 191] rest on the associative memory view.
@@
Multi-hop question-answering. A line of work has looked at how pretrained models respond to two-hop questions in natural language e.g., “What is the calling code of the birthplace of Frida Kahlo?” (example from Press et al. [135]). Results here have been limited [135, 21] or mixed [185, 186, 16]. Success has been shown on synthetic knowledge provided there is exponential amounts of data, or a curriculum, or very long amounts of training [188, 174]. Orthogonally, Wang et al. [178] identify that such in-weights implicit reasoning can be hurt by scaling up the parameters, perhaps due to “excessive memorization”, which we interpret as associative memorization. It is possible that some of these negative results may be attributed to the reversal curse [18, 9], or the lack of extended computation e.g., we use pause tokens [57]. But a dedicated study of this gap between our synthetic settings and these settings is important and left for future work.
Reversal curse and (a)symmetric knowledge. The reversal curse [18, 9] is a well-known out-of-distribution, in-weights failure mode of next-token-trained Transformers. Such models are unable to recall $u$ given $v$, when trained to recall $v$ given $u$, suggesting asymmetric storage in parametric memory. Fixes for the reversal curse have involved reversed or permuted data augmentation [64, 102, 88, 55]. In our settings, we find that reverse edges are critical to elicit implicit reasoning in our large path-star tasks (see §D.1), but is not necessary to elicit a geometry in the smaller graphs. [^13] It appears that the observations on small graphs in Khona et al. [85] do not require memorizing the reverse edges, which aligns with our findings on small graphs. Various theories have been proposed to understand this failure [97, 202, 173], which may be worth revisiting under a geometric view. One may also view our contrast between associative and geometric memory (of a generic graph data) as a generalization of the aforementioned contrast between the asymmetric and symmetric knowledge storage (of a more specific, disjoint set of associations). Future work may discover a nuanced connection between our work and the reversal curse.

### v3

@@ 4.3 Where does the low-rank spectral bias come from?
Open Question 3: Characterizing Associative-vs-Geometric Bias of Deeper Models (i) To what precise extent does a deep sequence model memorize different parts of a graph dataset geometrically vs. associatively? (ii) What local or global notions of graph complexity—such as the graph spectrum, connectivity, node & vertex count and degree—and what aspects of optimization—such as the architecture, learning rate, training time —control this bias? (iii) What interventions in the data, architecture or optimization can be made to make memorization more geometric or more associative as desired in practice?
In §C.3.1, we provide some preliminary analysis for (iii), by demonstrating how various optimization hyperparameters (e.g., learning rate, weight decay, initialization) affect this bias. We leave a dedicated analysis for future work.
4.4 Zigzagging and bipartite memory
Sometimes, we observe in various settings a zigzagging geometry: along certain top directions, adjacent vertices get embedded far away, while vertices that are exactly two-hop away get embedded close. This happens noticeably in two settings: (a) multi-layered models (especially when there is no weight decay) and/or (b) models with untied (un)embedding weights (e.g., Fig.˜9). We explain why zigzagging occurs, and clarify that this still falls within our extended definition of a geometric memory in §2.5. In short, zigzagging corresponds to negative eigendirections in the adjacency matrix, and these directions are picked up only when the architecture has the expressive power for it.
Zigzagging directions and negative eigenvalues. A key insight is that zigzagging directions correspond to negative eigendirections in the adjacency matrix. An adjacency matrix (without self-edges) has both positive and negative eigenvalues. Typically, the most negative/bottommost eigendirections assign, alternatingly, +1+1 and 1-1 values as we walk along a path. Likewise, if we were to consider a singular value decomposition (SVD) of the adjacency matrix—which becomes relevant for weight-untied models—such negative eigendirections are “folded up”: all strong eigendirections, with the most positive and most negative eigenvalues $\lambda$, become top singular directions with singular value $|\lambda|$ (see [209] for relevant facts on spectral graph theory). Thus, in all these cases, even though the geometry looks unruly, and even though the cosine heatmaps don’t reveal a multi-hop structure, the graph data has been stored by the model in a factorized form.
As discussed in §2.5, this storage is still geometric. Furthermore, the well-behaved “global” directions may still exist, and extractable via an appropriate probe on the embeddings. One way to test this is to add self-edges to the graph and see if the zigzags disappear from the top directions (as this would reduce the singular value of the zigzagging direction). We find this to be the case in §C.3.1 Fig.˜25. This also offers an explanation for recent observations that adding identity statements improves two-hop reasoning [104].
Why are zigzagging directions learned? We claim that such directions are picked up by the model (a) because they help drive the loss down and (b) because they can be expressed only by certain models. A model relying only on the top positive eigendirections would recover a positive definite approximation of the adjacency matrix, whose diagonals are highly positive i.e., the logit $f(u)[u]$ would be high since ${\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(u)$ would be high. Such a model would suffer a high softmax loss if there are no self-edges in the dataset. The negative eigendirections ${\bf{e}}_{\tt neg}$ help correct for these highly positive diagonal logits by introducing a negative definite component of the form ${\bf{e}}_{\tt neg}\lambda_{\tt neg}{\bf{e}}_{\tt neg}^{T}$ where $\lambda_{\tt neg}<0$. Indeed, when we add self-edges to the dataset, the zigzagging usually disappears!
Such zigzagging however can only be expressed by some architectures. A weight-tied Node2Vec model cannot, for such a model can only express positive definite matrices. A weight-untied model can express such directions as $({\bf{e}}_{neg})(-{\bf{e}}_{neg})^{T}$; indeed, weight-untied two-layer models are known to learn an SVD decomposition [211; 89; 162] which must include such directions as top singular directions. A multi-layered weight-tied model can also express these directions: it can use the intermediate layer to express the negative $\lambda_{\tt neg}$. Thus, we see zigzagging only where it is helpful (when no self-edges exist) and only when the model can express it (weight-untied or multi-layered).
@@ 4.4 Zigzagging and bipartite memory
Open question. Although visually unruly, these zigzagging eigendirections are highly-structured: in spectral graph theory, they are understood to be an approximate bipartite cut of the graph. Thus, they correspond to a form of geometric memory distinct from the “global” one, best termed as a bipartite geometric memory. Next, while the “global” geometry endows the model with navigational powers (as witnessed in §2), what benefits does the bipartite memory bring to a model’s reasoning abilities?
(a) Cycle
Figure 9: Embeddings of a weight-untied Transformer shows a zigzagging geometry. Both the 3D visualizations of the embeddings and their cosine similarities show a zigzagging pattern (adjacent vertices have very low, even negative, similarity).
@@ 5.1 Analysis of Transformer memory
Associative memory. The concept of associative memory dates back to theories of how information is stored in the brain [110; 193]. These ideas have since been explicitly modeled through various architectures such as Hopfield networks [76; 153], energy-based models [95], and other modern Transformer-style inventions [98; 75]. Closest to us are works that analyze how architectures implicitly behave as associative memory storage, such as in autoencoders [150] or Transformers [22; 28; 133; 54; 164; 174]. Others have established connections between the attention mechanism and Hopfield networks [153; 171]. We emphasize that this view is sufficient to understand Transformer behavior on disjoint facts, evidenced by the rich empirical literature built on this view. The geometric view only seems necessary when the facts become interdependent.
@@
Geometric memory as node representations. Although geometric memory has not been explicitly labeled as a distinct form of parametric memory, we find scattered observations of it in literature in weaker forms or more limited settings. [84; 91; 200; 134] show that nodes in graph-like datasets are embedded in a geometric way that encodes distances. These settings, which are on smaller-scale graphs, involve some form of global supervision or latent high-level structure, while we find that such a geometry can arise even with atomic facts. Jiang et al. [84] further juxtapose these structured embeddings against associative memory. However, they suggest that the structure becomes relevant only in the underparameterized setting, and that an associative memory suffices as an abstraction in the overparameterized setting; we find this is not the case. Independently, on two-hop datasets, [200; 184] report how models first “memorize” the training data, before “generalizing” on their reasoning task; under the hood, this translates to associative and geometric memorization respectively. Nishi et al. [134] argue and report that knowledge editing becomes tricky when memorizing cyclic graphs. Indeed, we suspect that a geometric memory poses problems to knowledge editing, since information is not stored loosely. This is an important open problem for future work to address.
@@
Geometric memory as factorized knowledge. Recall that the geometric memory can be viewed as storing a low-rank factorization of the adjacency matrix. The advantages of such factorizations is again scattered in literature. Saxe et al. [163, 162] prove that a two-layer model naturally learns an SVD factorization of input-output associations. This is a powerful result demonstrating the value of depth in learning hierarchical representations; but the global, multi-hop nature of these representations was not highlighted. Also, unlike our setting, theirs requires explicit pressures to demonstrate a low rank spectral bias (either through a bottleneck or early-stopping) for they study the squared-error loss. [^9] This requires some nuance. Under their squared error loss (applied without a softmax layer), their (weight-untied) fully-trained 2-layered model would learn the 1-hop adjacency matrix perfectly, with no global, multi-hop information revealed in the logits. Although the associations are stored in a factorized form, the final hidden representations do not skew further towards the higher directions than the eigenvalues themselves. For the cross-entropy loss, there is such a skew (Fig. 7). More recently, Huang et al. [80] point out that two-hop reasoning emerges if and only if the key-query matrices in an attention layer are kept separate (as against being collapsed into a single matrix); we show that such favorable effects occur in more general architectures without attention layers. They also highlight the role of an implicit regularization effect from their reasoning supervision; while this may indeed explain the benefits of a reasoning objective, we emphasize that this does not resolve why a geometry arises under pure local supervision—see Remark˜1.
@@
We also interpret these prior analyses [163; 80] as a study of how $2$-layered models memorize a special type of graphs: bipartite graphs (although they do not express it as such). [^10] Concretely, Saxe et al. [163] study mappings between “entity” tokens and binary feature vectors; we translate this as a bipartite graph where “entity” nodes are connected to various “feature” nodes. [80] study a mapping between “subject” nodes and “answer” nodes. Importantly, as noted in our own analysis in §4, such 2-layered models do not capture the competition between associative and geometric memory. [^11] At least not in its full glory; in $2$-layered weight-tied models, there is only a contrived form of associative memory (Proposition 5); if weight-untied, a natural form can arise in a single step (Proposition 4). For that, a study of deeper models is required, where the dynamics along the various spectral components become entangled; to keep them untangled, unnatural assumptions about the initializations need to be made, as shown in [162]. Another nuance in these prior 2-layered analyses is the lack of weight-tying, which creates significant differences in what is learned, as we identify in §4.4.
@@
(Implicit regularization in matrix factorization is an insufficient explanation of geometric memory) Huang et al. [80] attribute the success of reasoning to implicit regularization effects that arise in matrix completion tasks; we argue that these effects do not explain why a geometry arises over associative storage in our experiments. The “matrix completion” view in Huang et al. [80] is that during training the model attempts to fit certain observed entries in a $2$-hop adjacency matrix ${\bf{A}}^{2}$; any test query in effect demands that the model “complete” a missing $2$-hop entry. While many minima exist in such underdetermined matrix completion tasks, it is known [68; 12] that gradient descent on a deep factorized model finds the least nuclear norm matrix consistent with the observed entries. However, our edge-memorization-only task is a well-determined task that requires storing the adjacency matrix $\bf{A}$} as is, without any missing entries to infer; thus no such implicit pressure must exist. Yet a geometry arises in this task as shown in Section˜3.1. While it is still possible that there is a certain nuclear norm minimization effect even in edge-memorization tasks—perhaps because the logits themselves are not well-determined—the connection is not as obvious as it may seem.
@@
Expressive capacity. Theoretical works have quantified bounds on the expressive capacity of models when it comes to memorizing sequences [115; 87; 114; 86; 93] as opposed to associations between pairs of bigrams. These do not comment on the learning dynamics. These works typically assume that the token embeddings are all well-separated from each other (an assumption that empirically breaks in our setting). In light of the geometric view, it is necessary to restate expressive capacity bounds in terms of geometric capacity of a network and the “geometric complexity” of the dataset.
@@
Empirical analyses. Other works [10; 124; 155; 138] have performed careful empirical analyses of scaling laws for memorization, and quantified memorization in terms of “bits per parameter count”, known as bit complexity [181], which is related to our notion of bit count in Section˜3.2. All these analyses are worth revisiting with the newfound lens of a geometric memory. Zucchet et al. [216] empirically analyze the dynamics behind how facts are memorized in a model. Others [107; 208] have proposed methodological improvements to acquiring knowledge in a Transformer; of relevance to us is the finding in Zhang et al. [208] that training a model simultaneously on both facts and question-answering is a better way to integrate knowledge into the parameters.
@@
Mechanistic interpretability. There have been mechanistic investigations into how Transformers perform fact recall [113; 55] and where facts are stored in a transformer [54] and how it can be edited [214; 120]. Similar attempts have been made in traditional classifier networks [16; 116; 173]. Directly related to us are the works of Khona et al. [91] and Yao et al. [199]; Biran et al. [23] who perform a mechanistic interpretability analysis of how multi-hop recall works.
@@ 5.2 In-weights reasoning tasks
Synthetic graph tasks. Our work consolidates positive results of in-weights reasoning in literature, and presents a stronger instance of it, removing various confounders that make composition-learning easy. Khona et al. [91] report successful path-finding on $200$ nodes-large in-weights graphs, with varying path lengths and test-train overlap. Others [184; 47; 200; 80] report positive results on much shorter 2-hop tasks over a thousand entities or fewer. Geerts et al. [53] look at in-weights transitive inference, a special type of $\ell$-fold composition query where the model is trained on local comparisons and is queried on more distant comparisons. On settings with $7$ objects, Geerts et al. [53] find a clear difference between an in-context version of their task (where the model struggles) and an in-weights version (where the model succeeds). Our task of finding the first node is a much harder search task akin to finding the smallest node in an ordered relationship. Nagarajan et al. [128] discuss the limitations of next-token prediction, including on open-ended in-weights tasks, in lower data regimes. The fact that their next-token predictor achieves non-trivial performance on their in-weights task could be attributed to the effects of a geometric memory. Tangentially related is the positive finding in Yin and Wang [202] that the Transformer can compose in-weights knowledge given in-context demonstrations. It is worth noting that the (theoretical) arguments in both these works [128; 202] rest on the associative memory view.
@@
Multi-hop question-answering. A line of work has looked at how pretrained models respond to two-hop questions in natural language e.g., “What is the calling code of the birthplace of Frida Kahlo?” (example from Press et al. [146]). Results here have been limited [146; 23] or mixed [196; 197; 17]. Success has been shown on synthetic knowledge provided there is exponential amounts of data, or a curriculum, or very long amounts of training [199; 184]. Orthogonally, Wang et al. [188] identify that such in-weights implicit reasoning can be hurt by scaling up the parameters, perhaps due to “excessive memorization”, which we interpret as associative memorization. It is possible that some of these negative results may be attributed to the reversal curse [20; 9], or the lack of extended computation e.g., we use pause tokens [62]. But a dedicated study of this gap between our synthetic settings and these settings is important and left for future work.
Reversal curse and (a)symmetric knowledge. The reversal curse [20; 9] is a well-known out-of-distribution, in-weights failure mode of next-token-trained Transformers. Such models are unable to recall $u$ given $v$, when trained to recall $v$ given $u$, suggesting asymmetric storage in parametric memory. Fixes for the reversal curse have involved reversed or permuted data augmentation [70; 111; 94; 60]. In our settings, we find that reverse edges are critical to elicit implicit reasoning in our large path-star tasks (see §D.1), but is not necessary to elicit a geometry in the smaller graphs. [^12] It appears that the observations on small graphs in Khona et al. [91] do not require memorizing the reverse edges, which aligns with our findings on small graphs. Various theories have been proposed to understand this failure [105; 215; 183], which may be worth revisiting under a geometric view. One may also view our contrast between associative and geometric memory (of a generic graph data) as a generalization of the aforementioned contrast between the asymmetric and symmetric knowledge storage (of a more specific, disjoint set of associations). Future work may discover a nuanced connection between our work and the reversal curse.